OpenStudy (anonymous):

The sum of two numbers is 25 and their product is a maximum what are the two numbers?

OpenStudy (anonymous):

i would say \(12.5\)

OpenStudy (raden):

yeah, they are be fifty - fifty a= 25/2 b = 25/2 those are the numbers

OpenStudy (anonymous):

Could you explain yourself sat?

OpenStudy (anonymous):

An answer without an explanation doesnt really excite me a little

OpenStudy (anonymous):

there are many explanations the one i like best is by symmetry if you call one number \(x\) and the other one \(y\) with \(x+y=25\) then you cannot tell \(x\) and \(y\) apart, i.e. \(x+y=y+x\) since they are indistinguishable, it is impossible that one can be different from the other a most math teachery method it so call one \(x\) and the other one \(25-x\) and then maximize the product \(x(25-x)=25x-x^2\) max occurs at the vertex, \(-\frac{b}{2a}\) namely \(\frac{25}{2}\) same result is achieved via calculus